A full journal bearing of 50 mm diameter and 100 mm long has a bearing pressure of 1.4 N/mm2. The speed of the journal is 900 r.p.m. and the ratio of journal diameter to the diametral clearance is 1000. The bearing is lubricated with oil whose absolute viscosity at the operating temperature of 75°C may be taken as 0.011 kg/m-s. The room temperature is 35°C. Find : 1. The amount of artificial cooling required, and 2. The mass of the lubricating oil required, if the difference between the outlet and inlet temperature of the oil is 10°C. Take specific heat of the oil as 1850 J / kg / °C.
1
Expert's answer
2020-06-05T05:37:06-0400

As per the coefficient friction

we knows that


"\mu = \frac {33\times ZN \times d}{10^8 \times p \times c}"


on putting the value in above formula we get the value of "\mu= 0.004"


Now for load we know that,


"W= pdl = 1.4 \times 50 \times 100= 7000" N and the rubbing velocity willl be


"V= \frac{\pi \times D \times N }{60}" = 2.4 m/s


Artificial cooling required = Heat generated - Heat dissipated = "CA(t_b-t_a) - Cld (t_b-t_a)=71.5-28 = 43.5"


(ii) Mass required

"Q=ms\Delta T"


18500 = m "\times 1850 \times" 10


m=0.23 kg /min